# Sufficient digits and density estimation: A Bayesian nonparametric approach using generalized finite Pólya trees

### v1

@@ 4.1 Finite Pólya tree distributions
is a stochastic process with each $Y_{x_{1:n-1},0}\sim\mathrm{B}(\alpha_{x_{1:n-1},0},\alpha_{x_{1:j-1},1})$ and $Y_{x_{1:j-1},1}=1-Y_{x_{1:j-1},0}$, and where the $Y_{x_{1:n-1},0}$’s with $n\in\mathbb{N}$ are independent. Then $Y$ is said to be an infinite beta-distributed sequence with parameter $\alpha$, and for any $n\in\mathbb{N}$, $Y^{(n)}$ to be a finite beta-distributed sequence with parameter $\alpha^{(n)}$ and $Y^{(>n)}$ to be an infinite beta-distributed sequence with parameter $\alpha^{(>n)}$. For short, write
@@
The finite beta-distributed sequence induces a random PDF
@@
The random PDF $f\big{(}\cdot\,|\,Y^{(n)}\big{)}$ is constant on every interval $I_{x_{1:n}}$ with $x_{1:n}\in\{0,1\}^{n}$. Using $\Pi^{(n)}$ as a prior distribution generates a distribution of $X=.X_{1}X_{2}\ldots$ where $X\,|\,Y^{(n)}\sim f(\dot{\,}|\,Y^{(n)})$. Then by (5), for $j=1,\ldots,n$, conditioned on both $Y^{(n)}$ and $X_{1:j-1}$, the probability that $X_{j}=0$ is $Y_{X_{1:j-1},0}$, and .$X_{n+1}X_{n+2}\ldots\sim\mathrm{Unif}[0,1)$ is independent of $X_{1:n}$. Recall that $\alpha_{x_{1:j-1},x_{j}}$ is allowed to be 0: If $\alpha_{x_{1:j-1},x_{j}}=0$ then $Y_{x_{1:j-1},x_{j}}=x_{j}$ (see Section 2) which implies $X_{j}=1-x_{j}$
@@
Letting $n\to\infty$ in Definition 4.1 , we recover the definition of a Pólya tree $\Pi^{(\infty)}=\mathrm{PT}(\alpha,I)$ as given by Lavine [25]. We refer to $\Pi^{(\infty)}$ as the PT prior.
@@
Arguments for considering a finite Pólya tree distribution can be found in [26, 38]. For instance, $\Pi^{(\infty)}$ is not always the distribution of a random PDF unless the $\alpha$’s increase sufficiently rapidly which leads to problems of robustness, cf. [26]. Further, if $\Pi^{(\infty)}$ is the distribution of a random PDF, this random PDF is almost surely discontinuous almost everywhere, cf. [12].
@@ 4.2 Generalized finite Pólya tree distributions
for its distribution which is called a generalized finite Pólya tree of type 1 with parameter $(\alpha,I)$. We refer to $f\big{(}\cdot\,|\,Y^{(N)}\big{)}$ as the random GFPT1 PDF and to $\Pi_{1}$ as the GFTP1 prior. We also refer to the distribution of $(Y,N)$ as the GFTP1 prior (it will be clear by the context what is meant).
@@
.$X_{N+1}X_{N+2}\ldots\sim\mathrm{Unif}[0,1)$ (8)
@@
[…]
@@
the inferential goal is to derive the posterior distribution of $(Y,N)$. The posterior distribution of the random GFTP1 PDF is denoted $\Pi_{1}(\cdot\,|\,Z_{1:m})$ and called the posterior GFTP1 PDF:
@@
Our objectives in this paper are the posterior GFTP1 PDF and the posterior distribution of $N$ under the GFPT1 prior.
@@
for its distribution which is called a generalized finite Pólya tree of type 2 with parameter $(\alpha,\beta)$. We refer to $f\big{(}\cdot\,|\,Y^{(N)},R^{(N)}\big{)}$ as the random GFPT2 PDF and to $\Pi_{2}$ as the GFTP2 prior. We also refer to the distribution of $(Y,R,N)$ as the GFTP2 prior (it will be clear by the context what is meant).
@@
Let $X\sim\Pi_{2}\sim\mathrm{GFPT2}(\alpha,\beta)$. Conditioned on $(Y,R,N)$ (or just $(Y^{(N)},R^{(N)})$, .$X_{N+1}X_{N+2}\ldots\sim\mathrm{Unif}[0,1)$ is independent of $X_{1:N}$ which follows the random PMF given by the numerator in (13). The law of $X$ can be expressed by a Bayesian hierarchical model:
@@
[…]
@@
the inferential goal is to derive the posterior distribution of $(Y,R,N)$. The posterior distribution of the GFTP2 PDF is denoted $\Pi_{2}(\cdot\,|\,Z_{1:m})$ and called the posterior GFTP2 PDF:
@@
In this paper, we focus on the posterior GFTP2 PDF and the posterior distribution of $N$ under the GFPT2 prior, although the posterior distribution of $R$ may possibly be of interest as well.
@@ 4.3 Specification of parameters
It remains to specify the parameters $\alpha$ and $\beta$ of the PT, GFTP1, and GFTP2 priors introduced in Sections 4.1 and 4.2, and to specify a prior distribution of $N$. This section introduces some specific choices of $\alpha$ and $\beta$, which are later used in Section 7. Priors for $N$ are discussed later in Remark 16 and in Section 7.
@@
for some $\alpha_{0}>0$, then $\Pi^{(\infty)}$ is almost surely absolutely continuous, cf. [25].
Another choice is inspired by the extended Newcomb-Benford law (the general significant-digit law of [22]): Recall that in the present paper we consider binary number representations, cf. Section 1, but the case of the usual binary numeral system extends to the general case of base-$q$ number representations with $q\geq 2$ an integer, so to make this point clear we write $q$ instead of 2 (later in Sections 6 and 7.3 we consider results for the most common cases $q=2$ and $q=10$). Consider $X=.X_{1}X_{2}\ldots=\sum_{i=1}^{\infty}X_{i}q^{-i}$, where $q^{-i}$ is the interval length of the $i$th partition of $[0,1)$. Extend $X$ by considering a continuous random variable $Z>0$ with base-$q$ number representation $Z=q^{M+1}X$, where $M$ is its order of magnitude and $X_{1}\not=0$ is its leading digit. Suppose $Z$ satisfies the extended Newcomb-Benford law (in base-$q$), which means that $\log_{q}(qX)$ is uniformly distributed between 0 and 1 – one says that $Z$ spans all orders of magnitude. Equivalently, for all $n\in\mathbb{N}$ and $x_{1:n}\in\{0,\ldots,q-1\}^{n}$,
@@
Here, $(c_{n})_{n\geq 1}$ is a sequence of user-specified positive parameters that control the variance of $Y$. Specifically, $\mathrm{Var}(Y_{x_{1:n}})$ is inversely proportional to $c_{n}$. Finally, to complete the description of the distribution of $Z$, we need to specify a joint distribution of $M$ and $X$. We defer this to Sections 6 and 7.3.
@@ 5 Bayesian analysis
[…]
We use the following notation. Let the data be given by a a realization $z_{1:m}\in[0,1)^{m}$ of $Z_{1:m}$. For $j\in\mathbb{N}_{0}$ and $x_{1:j}\in\{0,1\}^{j}$, let $n_{x_{1:j}}(z_{1:m})$ be the number of observations $z_{i}$ ($i=1,\ldots,m$) falling in the interval $I_{x_{1:j}}$. By Definitions 4.4 and 4.5, we have almost surely that $n_{x_{1:j}}(Z_{1:m})=0$ if $\alpha_{x_{1:j}}=0$ and $j>1$, so assume $n_{x_{1:j}}(z_{1:m})=0$ if $\alpha_{x_{1:j}}=0$ and $j>1$.
@@ 5.1.1 Posterior simulation under a GFPT1 prior
We start by considering the posterior distribution of $(N,Y)$ for the case (10) where a GFTP1 prior has been specified. Let $\gamma(z_{1:m},\alpha)=(\gamma_{x_{1:n}}(z_{1:m},\alpha)\,|\,x_{1:j}\in\{0,1\}^{j},\,j\in\mathbb{N})$ be the infinite sequence with entries
@@
$p(n\,|\,z_{1:m})\propto p_{n}\frac{\prod_{j=1}^{n}\prod_{x_{1:j-1}\in\{0,1\}^{j-1}}\mathrm{b}\left(\gamma_{x_{1:j-1},0}(z_{1:m},\alpha),\gamma_{x_{1:j-1},1}(z_{1:m},\alpha)\right)}{{\prod_{x_{1:n}\in\{0,1\}^{n}}\ell_{x_{1:n}}^{n_{x_{1:n}}(z_{1:m})}}}.$ (20)
@@
The PMF in (20) does not belong to a known parametric family. Since $Y^{(N)}$ is of varying dimension $12^{N}-$ 1, one may suggest to use the reversible jump Markov chain Monte Carlo algorithm [13, 17] for simulating from the posterior density $p(y^{(n)}\,|\,z_{1:m})\propto p(y^{(n)},z_{1:m})$, cf. (B1). However, it is much easier to use Theorem 5.1, since $Y$ conditioned on $(N,Z_{1:m})$ has a conjugate BS-prior, whilst posterior simulation of $N$ is straightforward: First, $N\,|\,Z_{1:m}=z_{1:m}$ can be sampled either exactly, or approximately, or in an asymptotically exact way via a Metropolis-Hastings algorithm as follows. If $N_{\max}$ is an upper bound for $N$, then sample $N$ exactly from a categorical distribution over $\{0,\ldots,N_{\max}\}$ with unnormalized weights given by the right hand side in (20) for $n=0,\ldots,N_{\max}$. This method may also be used to provide an approximate simulation if $N_{\max}$ is a user-defined suitable upper bound for $N$. Alternatively, use a Metropolis-Hastings algorithm with equilibrium density given by (20) (we never found a need for such an algorithm in the examples considered later in this paper). Second, simply use (21) and (22) when simulating $Y$ conditioned on $N$ and $Z_{1:m}=z_{1:m}$.
The posterior mean of the random PDF in (6) is the optimal Bayesian point estimator under the squared-error loss function [34]. The following corollary becomes useful for the calculation of this estimator.
@@
The right hand side in (23) does not depend on $Y^{(N)}$ and the term $p(n\,|\,z_{1:m})$ may be calculated/approximated as discussed in Remark 11. Hence, the posterior mean in (23) may be determined either by a numerical approximation or by the Monte Carlo estimate of

### v2

@@ 4.1 Finite Pólya tree distributions
is a stochastic process with each $Y_{x_{1:n-1},0}\sim\mathrm{Beta}(\alpha_{x_{1:n-1},0},\alpha_{x_{1:j-1},1})$ and $Y_{x_{1:j-1},1}=1-Y_{x_{1:j-1},0}$, and where the $Y_{x_{1:n-1},0}$’s with $n\in\mathbb{N}$ are independent. Then $Y$ is said to be an infinite beta-distributed sequence with parameter $\alpha$, and for any $n\in\mathbb{N}$, $Y^{(n)}$ to be a finite beta-distributed sequence with parameter $\alpha^{(n)}$ and $Y^{(>n)}$ to be an infinite beta-distributed sequence with parameter $\alpha^{(>n)}$. For short, write
@@
The finite beta-distributed sequence induces a random PDF on $[0,1)$ which is given by
@@
The random PDF $f\big(\cdot\,|\,Y^{(n)}\big)$ is constant on every interval $I_{x_{1:n}}$ with $x_{1:n}\in\{0,1\}^{n}$. Using $\Pi^{(n)}$ as a prior distribution generates a distribution of $X=.X_{1}X_{2}\ldots$ where $X\,|\,Y^{(n)}\sim f(\dot{\,}|\,Y^{(n)})$. Then by (5), for $j=1,\ldots,n$, conditioned on both $Y^{(n)}$ and $X_{1:j-1}$, the probability that $X_{j}=0$ is $Y_{X_{1:j-1},0}$, and $X_{n+1}X_{n+2}\ldots\sim\mathrm{Unif}[0,1)$ is independent of $X_{1:n}$. Recall that $\alpha_{x_{1:j-1},x_{j}}$ is allowed to be 0: If $\alpha_{x_{1:j-1},x_{j}}=0$ then $Y_{x_{1:j-1},x_{j}}=x_{j}$ (see Section 2) which implies $X_{j}=1-x_{j}$ .
@@
Letting $n\to\infty$ in Definition 4.1 , we recover the definition of a Pólya tree $\Pi^{(\infty)}=\mathrm{PT}(\alpha,I)$ as given by Lavine (1992). We refer to $\Pi^{(\infty)}$ as the PT prior.
@@
Arguments for considering a finite Pólya tree distribution can be found in Lavine (1994) and Wong and Ma (2010). For instance, $\Pi^{(\infty)}$ is not always the distribution of a random PDF unless the $\alpha$’s increase sufficiently rapidly which leads to problems of robustness, cf. Lavine (1994). Further, if $\Pi^{(\infty)}$ is the distribution of a random PDF, this random PDF is almost surely discontinuous almost everywhere, cf. Ferguson (1974).
@@ 4.2 Generalized finite Pólya tree distributions
for its distribution which is called a generalized finite Pólya tree of type 1 with parameter $(\alpha,I)$. We refer to $f\big(\cdot\,|\,Y^{(N)}\big)$ as the random GFPT1 PDF and to $\Pi_{1}$ as the GFPT1 prior. We also refer to the distribution of $(Y,N)$ as the GFPT1 prior (it will be clear by the context what is meant).
For prior elictation pruposes, it might be of interest to understand the finite dimensional laws of $f(\cdot\,|\,Y^{(N)})$. The next proposition sheds light on some aspects of such finite dimensional laws by providing an explicit expression for the mean.
Proposition 4.5.
Let $P(\cdot)$ be the random probability measure given by the random PDF in (6). Using notation as in Definition 3.1, for every $d\in\mathbb{N}$ and $x_{1:d}\in\{0,1\}^{d}$, we have
$\mathrm{E}\big[P(I_{x_{1:d}})\big]=\sum_{n=0}^{d-1}p_{n}\,\frac{\ell_{x_{1:d}}}{\ell_{x_{1:n}}}\,\prod_{j=1}^{n}\frac{\alpha_{x_{1:j}}}{\alpha_{x_{1:j-1},0}+\alpha_{x_{1:j-1},1}}+\mathrm{P}(N\geq d)\prod_{j=1}^{d}\frac{\alpha_{x_{1:j}}}{\alpha_{x_{1:j-1},0}+\alpha_{x_{1:j-1},1}}.$
Along the same lines of Proposition 4.5 it is possible to obtain analytical expressions for higher order moments of $P(I_{x_{1:d}})$, including the variance. However, their expression is rather complex which makes them less useful for prior elicitation purposes. By the $\pi$-$\lambda$ theorem, the expectation of $P(\cdot)$ can be extended to all measurable sets. Observe that the expression in Proposition 4.5 reduces to the one in Lavine (1992) if $N=\infty$ almost surely.
@@
$X_{N+1}X_{N+2}\ldots\sim\mathrm{Unif}[0,1)$ (8)
@@
where $p=(p_{0},p_{1},\ldots)$ is the PMF of $N$. Here, $Y^{(N)}$ takes the interpretation of an unobserved/latent parameter to which a prior distribution is assigned in the latter two steps, so that $f\big(x\,|\,Y^{(N)}\big)$ follows a $\mathrm{GFPT1}(\alpha,I)$ prior. Moreover, $Y^{(>N)}\,|\,(X,Y^{(N)})\sim\mathrm{BS}(\alpha^{(>N)})$ depends only on $N$.
@@
the inferential goal is to derive the posterior distribution of $(Y,N)$. The posterior distribution of the random GFPT1 PDF is denoted $\Pi_{1}(\cdot\,|\,Z_{1:m})$ and called the posterior GFPT1 PDF:
@@
Our objectives in this paper are the posterior GFPT1 PDF and the posterior distribution of $N$ under the GFPT1 prior.
Remark 9.
The introduction of the latent variable $N$ is clearly motivated by the coupling result in Theorem 3.2. However, $N$ plays a crucial role also as a smoothing parameter that improves the fitting process even when it is not directly a parameter of the data generaing process. Specifically, the introduction of a prior on $N$ allows to learn a posteriori the smoothness required to represent the density balancing over-fitting (i.e., extremely wiggly density estimates which typically occur with standard PT posteriors) and under-fitting (i.e., density estimates that do not capture important nuisances in the data, which happens with finite PT posteriors if $n$ is too small).
@@
for its distribution which is called a generalized finite Pólya tree of type 2 with parameter $(\alpha,\beta)$. We refer to $f\big(\cdot\,|\,Y^{(N)},R^{(N)}\big)$ as the random GFPT2 PDF and to $\Pi_{2}$ as the GFPT2 prior. We also refer to the distribution of $(Y,R,N)$ as the GFPT2 prior (it will be clear by the context what is meant).
@@
Let $X\sim\Pi_{2}\sim\mathrm{GFPT2}(\alpha,\beta)$. Conditioned on $(Y,R,N)$ (or just on $(Y^{(N)},R^{(N)})$), $X_{N+1}X_{N+2}\ldots\sim\mathrm{Unif}[0,1)$ is independent of $X_{1:N}$ which follows the random PMF given by the numerator in (13). The law of $X$ can be expressed by a Bayesian hierarchical model:
@@
where $(Y^{(N)},R^{(N)})$ takes the interpretation of an unobserved/latent parameter to which a prior distribution is assigned in the latter two steps, so that $f\big(x\,|\,Y^{(N)},R^{(N)}\big)$ follows a $\mathrm{GFPT2}(\alpha,\beta)$ prior. Conditioned on $(X,Y^{(N)},R^{(N)})$ we have that $Y^{(>N)}\sim\mathrm{BS}(\alpha^{(>N)})$ and $R^{(>N)}\sim\mathrm{BS}(\beta^{(>N)})$ are independent and depend only on $N$.
@@
the inferential goal is to derive the posterior distribution of $(Y,R,N)$. The posterior distribution of the GFPT2 PDF is denoted $\Pi_{2}(\cdot\,|\,Z_{1:m})$ and called the posterior GFPT2 PDF:
@@
In this paper, we focus on the posterior GFPT2 PDF and the posterior distribution of $N$ under the GFPT2 prior, although the posterior distribution of $R$ may possibly be of interest as well.
@@ 4.3 Specification of parameters
It remains to specify the parameters $\alpha$ and $\beta$ of the PT, GFPT1, and GFPT2 priors introduced in Sections 4.1 and 4.2, and to specify a prior distribution of $N$. This section introduces some specific choices of $\alpha$ and $\beta$, which are later used in Section 6. Priors for $N$ are discussed later in Remark 17 and in Section 6.
@@
for some $\alpha_{0}>0$, then $\Pi^{(\infty)}$ is almost surely absolutely continuous, cf. Lavine (1992). Other choices include $\alpha_{x_{1:n}}=n^{\delta}$ or $\alpha_{x_{1:n}}=\delta^{n}$, see Watson et al. (2017) for further details. If $I$ is given by the standard dyadic partitions, an application of Proposition 4.5 entails that all such choices where $\alpha_{x_{1:n}}$ depends only on $n$ lead to a prior mean of the GFPT1 random density equal to the uniform density.
Another choice is inspired by the extended Newcomb-Benford law (the general significant-digit law of Hill (1995)): Recall that in the present paper we consider binary number representations, cf. Section 1, but the case of the usual binary numeral system extends to the general case of base-$q$ number representations with $q\geq 2$ an integer, so to make this point clear we write $q$ instead of 2 (later in Sections 4.4 and 6.3 we consider results for the most common cases $q=2$ and $q=10$). Consider $X=.X_{1}X_{2}\ldots=\sum_{i=1}^{\infty}X_{i}q^{-i}$, where $q^{-i}$ is the interval length of the $i$th partition of $[0,1)$. Extend $X$ by considering a continuous random variable $Z>0$ with base-$q$ number representation $Z=q^{M+1}X$, where $M$ is its order of magnitude and $X_{1}\not=0$ is its leading digit. Suppose $Z$ satisfies the extended Newcomb-Benford law (in base-$q$), which means that $\log_{q}(qX)$ is uniformly distributed between 0 and 1 – one says that $Z$ spans all orders of magnitude. Equivalently, for all $n\in\mathbb{N}$ and $x_{1:n}\in\{0,\ldots,q-1\}^{n}$,
@@
Here, $(c_{n})_{n\geq 1}$ is a sequence of user-specified positive parameters that control the variance of $Y$. Specifically, $\mathrm{Var}(Y_{x_{1:n}})$ is inversely proportional to $c_{n}$. Finally, to complete the description of the distribution of $Z$, we need to specify a joint distribution of $M$ and $X$. We defer this to Sections 4.4 and 6.3.
@@
4.4 Modelling multi-scale human data via Benford’s law and generalized Pólya trees
Real-world datasets – such as those involving incomes, city sizes, or seismic magnitudes – often span multiple orders of magnitude, see Clauset et al. (2009). Modelling such data poses a challenge for Bayesian density estimation because traditional methods may have difficulty capturing both the coarse-scale variability and the fine-scale structure (Tokdar and Cunningham, 2024). We show here how a simple extension of a GFPT1 prior can be used to obtain a simple, yet powerful, Bayesian model for density estimation in scenarios where data span multiple orders of magnitudes and their digits exhibit the Newcomb–Benford law. The main idea lies in independently modelling the order of magnitude of the data and their digits, by assuming a GFPT1 prior for the digits.
In the following, we write $\mathrm{BPT}_{q}$ for the $\mathrm{GFPT1}(I,\alpha)$ prior obtained when $I$ is the sequence of standard base-$q$ nested partitions of the unit interval (see Section 4.3) and $\alpha$ is as in (18). We refer to $\mathrm{BPT}_{q}$ as the base-$q$ Benford Pólya tree.
$\mathrm{P}(M_{i}=k\,|\,\omega)=\omega_{k},\quad k\in T,$
where we impose the prior
$\omega=(\omega_{k}\,|\,k\in T)\sim\mathrm{Dir}(\eta)$
The $\mathrm{MBPT}_{q}$ prior induces a scale invariant (in the terminology of Hill (1995)) random density on $T\times[0,1)$ equipped with the product of counting measure on $T$ and Lebesgue measure on $[0,1)$. This random density is given by
@@ 5 Bayesian analysis
This section deals with the inferential goals discussed in Remarks 8 and 10, considering a sample $Z_{1:m}$ of $m$ $[0,1)$-valued random variables with either a $\mathrm{GFPT1}$ or a $\mathrm{GFPT2}$ prior, cf. (10) and (14). The proofs of all theorems and one corollary in this section are found in the Supplemental material.
We use the following notation. Let the data be given by a a realization $z_{1:m}\in[0,1)^{m}$ of $Z_{1:m}$. For $j\in\mathbb{N}_{0}$ and $x_{1:j}\in\{0,1\}^{j}$, let $n_{x_{1:j}}(z_{1:m})$ be the number of observations $z_{i}$ ($i=1,\ldots,m$) falling in the interval $I_{x_{1:j}}$. By Definitions 4.4 and 4.6, we have almost surely that $n_{x_{1:j}}(Z_{1:m})=0$ if $\alpha_{x_{1:j}}=0$ and $j>1$, so assume $n_{x_{1:j}}(z_{1:m})=0$ if $\alpha_{x_{1:j}}=0$ and $j>1$.
@@ 5.1.1 Posterior simulation under a GFPT1 prior
We start by considering the posterior distribution of $(N,Y)$ for the case (10) where a GFPT1 prior has been specified. Let $\gamma(z_{1:m},\alpha)=(\gamma_{x_{1:n}}(z_{1:m},\alpha)\,|\,x_{1:j}\in\{0,1\}^{j},\,j\in\mathbb{N})$ be the infinite sequence with entries
@@
$p(n\,|\,z_{1:m})\propto p_{n}\frac{\prod_{j=1}^{n}\prod_{x_{1:j-1}\in\{0,1\}^{j-1}}\mathrm{beta}\left(\gamma_{x_{1:j-1},0}(z_{1:m},\alpha),\gamma_{x_{1:j-1},1}(z_{1:m},\alpha)\right)}{{\prod_{x_{1:n}\in\{0,1\}^{n}}\ell_{x_{1:n}}^{n_{x_{1:n}}(z_{1:m})}}}.$ (20)
@@
The PMF in (20) does not belong to a known parametric family. Since $Y^{(N)}$ is of varying dimension $2^{N}-1$, one may suggest to use the reversible jump Markov chain Monte Carlo algorithm (Geyer and Møller, 1994; Green, 1995) for simulating from the posterior density $p(y^{(n)}\,|\,z_{1:m})\propto p(y^{(n)},z_{1:m})$, cf. (B1). However, it is much easier to use Theorem 5.1, since $Y$ conditioned on $(N,Z_{1:m})$ has a conjugate BS-prior, whilst posterior simulation of $N$ is straightforward: First, $N\,|\,Z_{1:m}=z_{1:m}$ can be sampled either exactly, or approximately, or in an asymptotically exact way via a Metropolis-Hastings algorithm as follows. If $N_{\max}$ is an upper bound for $N$, then sample $N$ exactly from a categorical distribution over $\{0,\ldots,N_{\max}\}$ with unnormalized weights given by the right hand side in (20) for $n=0,\ldots,N_{\max}$. This method may also be used to provide an approximate simulation if $N_{\max}$ is a user-defined suitable upper bound for $N$. Alternatively, use a Metropolis-Hastings algorithm with equilibrium density given by (20) (we never found a need for such an algorithm in the examples considered later in this paper). Second, simply use (21) and (22) when simulating $Y$ conditioned on $N$ and $Z_{1:m}=z_{1:m}$.
The posterior mean of the random PDF in (6) is the optimal Bayesian point estimator under the squared-error loss function (Robert, 2007). The following corollary becomes useful for the calculation of this estimator.
@@
In the right hand side of (23), the term $p(n\,|\,z_{1:m})$ may be calculated as discussed in Remark 12. Hence, the posterior mean in (23) may be determined either by a numerical approximation or by the Monte Carlo estimate of